### v1

# generalised polynomial chaos

Abstract
This paper presents an approach for the modelling of dependent random variables using generalised polynomial chaos. This allows to write chance-constrained optimization problems with respect to a joint distribution modelling dependencies between different stochastic inputs. Arbitrary dependencies are modelled by using Gaussian copulas to construct the joint distribution. The paper exploits the problem structure and develops suitable transformations to ensure tractability. The proposed method is applied to a probabilistic power reserve procurement problem. The effectiveness of the method to capture dependencies is shown by comparing the approach with a standard approach considering independent random variables.
I Introduction
distributed energy resources are increasingly participating in ancillary service (AS) markets. Many of these resources. such as heat pumps, electric vehicles and photovoltaic (PV) installations have uncertain capabilities. To facilitate the participation of these DERs, some transmission system operators already allow bids for which the full power availability is only guaranteed with a certain probability [1]. This has important consequences for the reliable procurement of reserve capacity. Add short review on approaches for reserve procurement. Look for probabilistic approaches + especially considering dependencies of multiple DER types.
Recently, generalised polynomial chaos (gPC) has been used in power systems applications to model propagation of uncertain input variables. Advantages include the ability to model general probability distributions and represent stochastic variables through orthogonal polynomial expansions. This approach was first used for stochastic power flow computations [4], with inequality constraints formulated only with respect to the expected value. Later work, such as [5], introduces the use of gPC for the solution of optimal power flow problems. Typically, approaches modelling uncertainty with gPC formulate chance constraints through a moment-based reformulation. For Gaussian random variables or distributionally robust (DR) uncertainty, exact reformulations are found [6]. For other distributions, [3] proposes an iterative tuning approach for the bounds in the moment-based reformulation in an attempt to recover the desired constraint satisfaction rate. A major limitation of these approaches, is that they consider the different random input variables to be independent. To allow modelling dependencies between random variables, [7] proposes to construct an orthonormal polynomial basis using a Gram-Schmidt (GS) procedure. A challenge with this approach is that it leads to numerical issues due to the poor conditioning of the GS procedure[2]. Other methods to model dependent random variables with gPC exist, as reported for example in [2]. The Rosenblatt transformation is a mapping that transforms dependent random variables to a space of independent random variables, in which an orthonormal polynomial basis can be constructed using standard approaches. This has for example been applied to solve stochastic power flows [9]. However, this approach requires to determine the inverse Rosenblatt transform for each sample used to obtain polynomial chaos expansion (PCE) coefficients. here further works on dependent gPC?
None of the works above allow to assess the probability of violating the power reserve requirements when considering uncertain bids exhibiting dependencies. Additionally, the proposed approaches to extend gPC to dependent variables, lead to computational issues due to error propagation and computation time. Therefore, this paper makes the following contributions:
  1. 1.
    An approach to model arbitrary dependencies between random variables using (Gaussian) copulas.
  2. 2.
    A tractable computation scheme for the gPC basis, using quadrature integration and exploiting the copula structure.
  3. 3.
    A convex reformulation for probabilistic power reserve procurement considering dependent random variables.
II Methodology
II-A gPC with Dependent Random Variables
Consider the random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{T}$ with joint PDF $p(\bm{\xi})$ and support $\Xi$.
A known method to deal with uncertainty (is the use of a Polynomial Chaos Expansion (PCE), casting a function $f(\bm{\xi})$ of random variables $\bm{\xi}$ into a sum of orthonormal polynomials $\psi(\bm{\xi})$:
$f(\bm{\xi})=\sum_{l=0}^{\infty}a_{l}\psi_{l}(\bm{\xi}),$ (1)
where orthonormality is required with respect to the expectation value
$\mathbb{E}[\psi_{m}(\bm{\xi})\psi_{n}(\bm{\xi})]=\int_{\Xi}\psi_{m}(\bm{\xi})\psi_{n}(\bm{\xi})p(\bm{\xi})d\bm{\xi}=\delta_{mn}.$ (2)
The coefficients can be computed by projection onto the polynomial basis, where the scalar product is identical to the expectation value of the product:
$a_{l}=\left\langle f(\bm{\xi}),\psi_{l}(\bm{\xi})\right\rangle=\mathbb{E}[f(\bm{\xi})\psi_{l}(\bm{\xi})]$ (3)
The main difficulty is the search for orthonormal polynomials with respect to the arbitrary PDF $p(\xi)$. As proven in (XXXSomething), polynomials arising from Askey-scheme are orthogonal for their respective probability distribution (e.g. Hermite polynomials for Gaussian PDFs), however, this assumes independence between the elements of the random vector. In turn, Navarro et al. suggested in [7] that for arbitrarily correlated random variables, orthonormal polynomials can be constructed from a set of linearly independent polynomials using the Gram-Schmidt orthonormalization scheme.
II-B Joint Probability using Copulas
For practical applications, the joint PDF $p(\bm{\xi})$ needs to be approximated in a tractable manner. This requires both a model of the marginal PDFs of each random variable $\xi_{i}$ as well as their dependence. Dependence modelling can be achieved by using a copula $C$, a function mapping the cumulative distribution functions (CDF) $F_{i}(\xi_{i})$ of each marginal to a joint CDF $F(\bm{\xi})$:
$F(\bm{\xi})=C(F_{1}(\xi_{1}),\ldots,F_{d}(\xi_{d}))$ (4)
As the copula takes CDFs as inputs, it effectively decouples the marginal distributions from the dependence structure. According to Sklar’s theorem (XXX), there exists a unique copula for every possible joint CDF. In practice, one can use an empirical copula, as for example in (Pål paper?), which makes no assumptions on the correlation structure or fit a parametric copula. Archimedean copulas such as Gumbel or Clayton copulas, have a single parameter to capture upper and lower tail-dependence respectively. Gaussian, or more generally Student-t copulas, are defined by a correlation matrix $\Sigma$ and degrees of freedom $\nu$, where in the limit $\nu\rightarrow\infty$, the Student-t becomes a Gaussian copula.
For the application reported here, featuring a high dimensionality a Gaussian copula was chosen. This allows for scalability and mathematically convenient expressions (XXX and what?). Setting $F_{i}(\xi_{i})=u_{i}$, it is defined as:
$C_{\bm{\Sigma}}(u_{1},...,u_{d})=\Phi_{\bm{\Sigma}}(\Phi^{-1}(u_{1}),...,\Phi^{-1}(u_{d}))$ (5)
where $\Phi^{-1}(u_{i})$ is the inverse of the normal CDF and $\Phi_{\bm{\Sigma}}(\bm{z})$ the CDF of the multivariate normal with correlation matrix $\bm{\Sigma}$. The joint PDF of the random vector $\bm{\xi}$ is then expressed as
$p(\bm{\xi})=c(\bm{u})*\prod_{i=1}^{d}f_{i}(\xi_{i})$ (6)
where $c(\bm{u})$ is the copula density
$c(\bm{u})=\frac{\partial^{d}C(u_{1},\dots,u_{d})}{\partial u_{1}\,\partial u_{2}\cdots\partial u_{d}}.$ (7)
In the case of a Gaussian copula, this results in
$c_{\Sigma}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)},$ (8)
where $\phi_{\bm{\Sigma}}(\bm{x})$ is the PDF of the multivariate normal with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
II-C Quadrature Integration
Both for the computation of the orthonormal polynomials using GS and for the projection of expansion coefficients in Eq. (3), integrals of the form
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}$ (9)
have to be evaluated numerically, where $g(\bm{\xi})$ is an arbitrary function. Given the high dimensionality of the integral, evaluation with Monte Carlo sampling or importance sampling schemes would be standard, this however leads to relatively high errors. Further, importance sampling schemes are difficult to implement, as the integrands do not follow a constant distribution
To allow accurate evaluation, quadrature integration is used, this however suffers from computational tractability problems if applied to high-dimensional problems. For this reason a scheme reducing the integration dimension is introduced. First, the integration domain needs to be transformed to a space with known quadrature rules. One such space is the unit cube $0,1]^{d}$ over which the Gauss-Legendre rules can be applied. thus let
$u_{i}=F_{i}(\xi_{i})\in[0,1],\qquad i=1,\dots,d.$ (10)
The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$, hence
$d\bm{\xi}=\frac{d\bm{u}}{\prod_{i=1}^{d}f_{i}(\xi_{i})}.$ (11)
Applying the change of variables and substituting in the definition of $p(\bm{\xi})$ from Eq. (6) yields
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}f\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
where $\bm{F}^{-1}(\bm{u})$ is the componentwise inverse CDF. This expression makes no assumptions on the copula used and hence… However, as the copula density $c(\bm{u})$ is in general neither polynomial nor constant, the quadrature will be inexact. In the specific case of Gaussian copulas, a transformation to latent Gaussian space
$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty),\qquad i=1,\dots,d.$ (13)
is helpful. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, hence
$d\bm{u}=\Big(\prod_{i=1}^{d}\phi(z_{i})\Big)\,d\bm{z}.$ (14)
Applying the change of variables to the copula density and using the definition of the Gaussian copula density from Eq. (8) yields
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\Phi(\bm{z}))\big)\;\phi_{\Sigma}(\bm{z})\,d\bm{z}.$ (15)
This form allows for Gauss-Hermite quadrature, which expects a polynomial weighted by an uncorrelated normal distribution. Computing the Cholesky factorisation of the correlation matrix $\bm{\Sigma}=\bm{L}\bm{L^{T}}$ allows to “colour“ the correlation onto the quadrature. Letting $\bm{y}_{n}$ and $w_{n}$ be the quadrature nodes and weights. The linear transformation
$\bm{z_{n}}=\bm{L}\bm{y_{n}}$ (16)
induces the correlation. Overall,
$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})\,d\bm{x}\;\approx\;\sum_{n=1}^{N}w_{n}\cdot g(\bm{F}^{-1}(\Phi(\bm{Ly_{n}})))$ (17)
Standard Gauss-Hermite quadrature uses a tensor-product grid, meaning evaluating $k$ points per dimension for a $d_{I}$ dimensional integral leads to $N=k^{d_{I}}$ nodes. The exponential scaling with the integral dimension requires a strict limitation.
II-C1 Numerical Improvements
Two ideas allow for decent computational performance when computing the gPC basis. First, GS sequentially evaluates inner products of polynomials to construct the basis polynomials. As noted in [8], the inner product of two arbitrary polynomials $P_{m}(\bm{\xi}),P_{n}(\bm{\xi})$ constructed from the multivariate monomial set $\{1,\xi_{1},\ldots,\xi_{d}^{\nu}\}$, can be expanded to a sum of monomial expectation values
$\left\langle P_{m}(\bm{\xi}),P_{n}(\bm{\xi})\right\rangle$ $=p_{0}\left\langle 1,1\right\rangle+p_{1}\left\langle 1,\xi_{1}\right\rangle+\ldots+p_{P}\left\langle\xi_{d}^{\nu},\xi_{d}^{\nu}\right\rangle$
$=e_{0}+e_{1}\mathbb{E}[\xi_{1}]+\ldots+e_{E}\mathbb{E}[\xi_{d}^{2\nu}]$ (18)
where $p_{i}$ and $e_{i}$ are the appropriate coefficients. Evaluating the monomial expectation values before applying GS can be done in a parallel fashion, allowing for increased speed.
Second, the expectation values of monomials only depends on dimensions appearing in the monomial, for instance
$\mathbb{E}[\xi_{1}\xi_{2}]=\int_{\Xi_{1,2}}\xi_{1}\xi_{2}\,p_{1,2}(\xi_{1},\xi_{2})\,d\bm{\xi}_{1,2}.$ (19)
For elliptic copulas the reduced PDFs have identical structure to the full joint PDF, so Eq. (6) still holds, but with less input variables.
These two observations limit the maximum integral dimension to be evaluated to $2\nu$ with expectations of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$. Crucially, this decouples integral dimension $d_{I}$ from random variable dimension $d$, avoiding the curse of dimensionality.
III Use Case
IV Conclusion