### v2

# generalised polynomial chaos

Abstract
This paper presents an approach for the modelling of dependent random variables using generalised polynomial chaos. This allows to write chance-constrained optimization problems with respect to a joint distribution modelling dependencies between different stochastic inputs. Arbitrary dependencies are modelled by using Gaussian Copulas to construct the joint distribution. The paper exploits the problem structure and develops suitable transformations to ensure tractability. The proposed method is applied to a probabilistic power reserve procurement problem. The effectiveness of the method to capture dependencies is shown by comparing the approach with a standard approach considering independent random variables.
I Introduction
distributed energy resources are increasingly participating in ancillary service (AS) markets. Many of these resources. such as heat pumps, electric vehicles and photovoltaic (PV) installations have uncertain capabilities. To facilitate the participation of these DERs, some transmission system operators already allow bids for which the full power availability is only guaranteed with a certain probability [1]. This has important consequences for the reliable procurement of reserve capacity. This raises new questions with respect to the amount of reserve that should be procured, as the availability of the procured reserve becomes a stochastic quantity.
Recently, generalised polynomial chaos (gPC) has been used in power systems applications to model propagation of uncertain input variables. Advantages include the ability to model general probability distributions and represent stochastic variables through orthogonal polynomial expansions. This approach was first used for stochastic power flow computations [5], with inequality constraints formulated only with respect to the expected value. Later work, such as [6], introduces the use of gPC for the solution of optimal power flow problems. Typically, approaches modelling uncertainty with gPC formulate chance constraints through a moment-based reformulation. For Gaussian random variables or distributionally robust (DR) uncertainty, exact reformulations are found [7]. For other distributions, [3] proposes an iterative tuning approach for the bounds in the moment-based reformulation in an attempt to recover the desired constraint satisfaction rate. A major limitation of these approaches, is that they consider the different random input variables to be independent. To allow modelling dependencies between random variables, [8] proposes to construct an orthonormal polynomial basis using a Gram-Schmidt (GS) procedure. A challenge with this approach is that it leads to numerical issues due to the poor conditioning of the GS procedure[2]. Other methods to model dependent random variables with gPC exist, as reported for example in [2]. The Rosenblatt transformation is a mapping that transforms dependent random variables to a space of independent random variables, in which an orthonormal polynomial basis can be constructed using standard approaches. This has for example been applied to solve stochastic power flows [11]. However, this approach requires to determine the inverse Rosenblatt transform for each sample used to obtain polynomial chaos expansion (PCE) coefficients. For linear dependencies, this can be simplified to the Nataf transform [4], but the authors of [2] show that it performs worse than their proposed GS procedure. Other approaches, so-called domination methods, approximate functions of dependent random variables with functions of independent random variables, but only show good performance when an accurate approximation can be obtained [2].
None of the works above handling reserve procurement allow to assess the probability of violating the power reserve requirements when considering uncertain bids exhibiting dependencies. Additionally, the proposed approaches to extend gPC to dependent variables, lead to computational issues due to error propagation and computation time. Therefore, this paper makes the following contributions:
  1. 1.
    An approach to model arbitrary dependencies between random variables using (Gaussian) Copulas.
  2. 2.
    A tractable computation scheme for the gPC basis, using quadrature integration and exploiting the Copula structure.
  3. 3.
    A convex reformulation for probabilistic power reserve procurement considering dependent random variables.
II Methodology
II-A gPC with Dependent Random Variables
Consider the random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{\top}$ of dimension $d$ with joint probability density function (PDF) $p(\bm{\xi})$ and support $\Xi$. As shown in previous works (e.g. [8]), a square-integrable random function $q(\bm{\xi})$ can be represented by a PCE
$q(\bm{\xi})\cong\sum_{l=0}^{L}a_{l}\psi_{l}(\bm{\xi}).$ (1)
where the equality becomes exact as the arbitrary number of terms $L$ goes to infinity. The basis polynomials $\{\psi_{l}\}$ are orthonormal with respect to the inner product
$\left\langle\psi_{m},\psi_{n}\right\rangle=\int_{\Xi}\psi_{m}(\bm{\xi})\,\psi_{n}(\bm{\xi})\,p(\bm{\xi})\,\mathrm{d}\bm{\xi}.$ (2)
The expansion coefficients $\{a_{l}\}$ are obtained by projection onto the polynomial basis. Since this inner product corresponds to an expectation, the coefficients read
$a_{l}=\left\langle q,\psi_{l}\right\rangle=\mathbb{E}\!\left[q(\bm{\xi})\,\psi_{l}(\bm{\xi})\right].$ (3)
The main challenge lies in the construction of orthonormal polynomial bases with respect to an arbitrary joint PDF $p(\bm{\xi})$. For independent random variables, orthogonal polynomial families associated with the Askey scheme can be employed, since they are orthogonal with respect to their corresponding marginal distributions (e.g. Hermite polynomials for Gaussian random variables).
For general dependent random variables, the basis can be constructed from a suitably chosen set of multivariate monomials of the random variables (e.g. all monomials with degree $\leq\nu$). As noted, a GS procedure can be used for the construction, but is numerically ill-conditioned [2]. To address this, [10] provides an alternative algorithm based on a Whitening transformation, relying only in part on GS procedures.
Finally, the moments of the output distribution of $q(\bm{\xi})$ can be recovered from the expansion coefficients. For instance the mean and variance are given by
$\mu$ $=a_{0}$ (4)
$\sigma$ $\cong\left\|[a_{1},\ldots,a_{L}]^{\top}\right\|_{2}$
II-B Joint Probability using Copulas
For practical modelling of statistically dependent random variables, the joint PDF $p(\bm{\xi})$ must be approximated in a tractable manner. This requires specifying both the marginal distributions of the individual components $f(\xi_{i})$ as well as their dependence structure.
Copulas provide a general and flexible framework for separating the modelling of marginal distributions from the modelling of statistical dependence. Specifically, a Copula $C$ links the marginal cumulative distribution functions (CDFs) $F_{i}(\xi_{i})$ to the joint CDF $F(\bm{\xi})$. This link is implicitly expressed as
$F(\bm{\xi})=C\!\left(F_{1}(\xi_{1}),\ldots,F_{d}(\xi_{d})\right).$ (5)
Sklar’s Theorem states, that for every multivariate CDF there exists a unique Copula capturing the general dependence structure given by (5). In practice, one of several parametric Copula families is used.
For the application considered in this paper no specific tail-dependence is expected. As a result, a Gaussian Copula is selected to maintain computational tractability and leverage its closed-form mathematical expressions. Identifying $u_{i}=F_{i}(\xi_{i})$, it is defined as:
$C_{\bm{\Sigma}}(\bm{u})=\Phi_{\bm{\Sigma}}(\Phi^{-1}(u_{1}),...,\Phi^{-1}(u_{d}))$ (6)
where $\Phi_{\bm{\Sigma}}(\bm{z})$ is the multivariate normal CDF with correlation matrix $\bm{\Sigma}$ and $\Phi^{-1}(u_{i})$ the inverse normal CDF. The joint PDF is then expressed as
$p(\bm{\xi})=c(\bm{u})*\prod_{i=1}^{d}f_{i}(\xi_{i})$ (7)
where $c(\bm{u})$ is the Copula density. For a Gaussian Copula, the density is given by
$c_{\bm{\Sigma}}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)},$ (8)
where $\phi_{\bm{\Sigma}}(\bm{x})$ is the multivariate normal PDF with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
II-C Quadrature Integration
Both the construction of the orthonormal polynomial basis and the computation of the expansion coefficients require the evaluation of a large number of integrals of the form
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}$ (9)
with $g(\bm{\xi})$ an arbitrary function. In general, these integrals cannot be evaluated analytically and must be approximated numerically. Given the potentially high dimensionality of the integrals, Monte Carlo sampling is a natural choice. This however requires a large amount of samples to achieve sufficient accuracy and sampling errors may accumulate throughout the PCE process. Moreover, the use of importance sampling is not straightforward in this context, since the integrands encountered vary across polynomial orders and random functions.
Therefore, to enable fast and accurate numerical integration, quadrature integration schemes in combination with dimensionality-reduction strategies are used. First, the integration domain is mapped to a space with known quadrature rules. One such space is the unit cube $0,1]^{d}$ on which tensor-grid Gauss-Legendre quadrature can be applied. Let
$u_{i}=F_{i}(\xi_{i})\in[0,1],\qquad i=1,\dots,d.$ (10)
The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$ the marginal PDFs. Hence
$d\bm{\xi}=\frac{d\bm{u}}{\prod_{i=1}^{d}f_{i}(\xi_{i})}.$ (11)
Applying the change of variables and substituting $p(\bm{\xi})$ from (7) yields
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}g\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
where $\bm{F}^{-1}(\bm{u})$ is short-hand notation for the componentwise inverse marginal CDF. Evidently, this transformation is valid no matter the Copula used, in principle allowing any Copula family to be used for this quadrature integration scheme. However, the Copula density $c(\bm{u})$ is in general neither polynomial nor constant, meaning Gauss-Legendre quadrature will be inexact.
For the specific case of Gaussian Copulas, consider a second transformation to a Gaussian latent space
$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty),\qquad i=1,\dots,d$ (13)
on which tensor-grid Gauss-Hermite quadrature is applicable. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, hence
$d\bm{u}=\Big(\prod_{i=1}^{d}\phi(z_{i})\Big)\,d\bm{z}$ (14)
Applying the change of variables and substituting the Gaussian Copula density from (8) yields
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\bm{\Phi}(\bm{z}))\big)\;\phi_{\bm{\Sigma}}(\bm{z})\,d\bm{z}$ (15)
with $\bm{\Phi}(\bm{z})$ the componentwise normal CDF. The result in (15) features a correlated normal PDF. However, Gauss-Hermite quadrature expects to weighted by an uncorrelated normal PDF. To address this, let $\bm{\Sigma}=\bm{L}\bm{L^{T}}$ be the Cholesky factorisation of the correlation matrix. For a given set of quadrature nodes $\{\bm{z}_{n}\}$ and weights $\{w_{n}\}$ the linear transformation
$\bm{y}_{n}=\bm{L}\bm{z}_{n}$ (16)
allows to “colour“ the correlation onto the quadrature nodes resulting in
$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})\,d\bm{x}\;\approx\;\sum_{n=1}^{N}w_{n}\cdot g(\bm{F}^{-1}(\Phi(\bm{Lz}_{n})))$ (17)
The number of nodes is given by $N=k^{d_{I}}$ with $k$ the number of evaluation points per dimension and $d_{I}$ the integral dimensionality. Due to this exponential scaling, dimensionality reduction strategies are necessary. Note that while the second transformation eliminated the non-polynomial Copula density $c(\bm{u})$, improving numerical accuracy, there is still no exactness guarantee. Gauss-Hermite quadrature is exact for polynomials up to degree 2k12k-1 weighted by a normal PDF, but $g(\bm{F}^{-1}(\bm{\Phi}(\bm{Lz})))$ does not necessarily behave like a polynomial of degree $\leq 2k-1$ in Gaussian latent space.
II-C1 Numerical Improvements
The dimensionality-reduction strategy used to make quadrature integration tractable stems from two key observations: First, as noted in [10], inner products between polynomials can be expanded to a sum of monomial expectation values. For instance, let $P_{m}(\bm{\xi}),P_{n}(\bm{\xi})$ be two arbitrary polynomials constructed from the set of multivariate monomials of degree $\leq\nu$ $\{1,\xi_{1},\ldots,\xi_{d}^{\nu}\}$. Then the inner product
$\left\langle P_{m},P_{n}\right\rangle$ $=d_{0}\!\left\langle 1,1\right\rangle+d_{1}\!\left\langle 1,\xi_{1}\right\rangle+\ldots+d_{D}\!\left\langle\xi_{d}^{\nu},\xi_{d}^{\nu}\right\rangle$
$=e_{0}+e_{1}\mathbb{E}[\xi_{1}]+\ldots+e_{E}\mathbb{E}[\xi_{d}^{2\nu}]$ (18)
where $d_{i},e_{i}$ are the appropriate coefficients. Second, the expectation values of monomials only depend on dimensions appearing in the monomial, for instance
$\mathbb{E}[\xi_{1}\xi_{2}]=\int_{\Xi_{1,2}}\xi_{1}\xi_{2}\,p_{1,2}(\xi_{1},\xi_{2})\,d\bm{\xi}_{1,2}.$ (19)
where $p_{1,2}(\xi_{1},\xi_{2})$ is a reduced joint PDF. For elliptic Copulas they have identical structure to the full joint PDF defined in (7), just with less input variables. For other Copulas, the reduced joint PDFS might need to be retrieved separately.
Now, instead of directly applying an orthonormalization scheme, the monomial expectation values of all pairs of monomials in the chosen set can be evaluated beforehand. When applying the orthonormalization scheme, the inner products can be expanded as in (18) and the expectation values substituted.
Crucially, precomputing expectation values can be done in a parallel fashion and naturally limits the maximum integral dimension to be evaluated to $2\nu$. This because the highest dimensional expectation values to be evaluated are of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$. This approach decouples the integral dimension $d_{I}$ from the random variable dimension $d$.
In the case where multiple random functions $q_{i}(\bm{\xi_{i}})$ only depending on subsets $\bm{\xi_{i}}$ of the random vector are expanded to the same polynomial basis, a similar approach can be used for the expansion coefficients. Here the expectation values $\mathbb{E}[q_{i}(\bm{\xi_{i}})],\ldots,\mathbb{E}[q_{i}(\bm{\xi_{i}})\xi_{d}^{\nu}]$ can be computed before with reduced dimension. Note, that the reduction in this case is dependent on the dimensionality of the random functions $q_{i}(\bm{\xi_{i}})$.
Finally, the set of monomials from which the polynomial basis is constructed can be tuned. For instance, assume the random functions $q_{i}(\bm{\xi_{i}})$ are models of PV power output at locations $i$ with $\bm{\xi_{i}}$ the local weather conditions like irradiation and temperature. Physically meaningless monomials such as the product of squares of the irradiations at two different locations can be dropped from the set without significantly affecting the expansion accuracy.
III Use Case
Reserve power, procured from AS markets, is activated whenever the grid frequency exceeds or dips below certain thresholds to ensure stable frequency levels. Given the critical nature of reserve activation, the risk of under-delivery must be kept to a minimum. This need for security can lead to high procurement costs. A recent initiative to keep economic efficiency is the Allocation of Cross-zonal Capacity and Procurement of aFRR Cooperation Agreement (ALPACA) which co-optimizes the procurement of reserves across borders of Austria, Czechia and Germany.
To validate the formalism introduced in this paper, a simple cross-border joint procurement optimization with uncertain bids and chance constraints on the available power is solved.
Formally, let $X$ and $Y$ be two zones sharing a common market with a set of bids $E_{i}(\bm{\xi})$ and costs $c_{i}$. Bids $i={1,\ldots,N_{X}}$ are located in zone $X$ and $i={N_{X}+1,\ldots,N_{X}+N_{Y}}$ are located zone Y. The decision variables $x_{i},y_{i}\in[0,1]$ determine the procured share of each bid. The objective is to minimize total cost, while both zones meet their reserve needs $R^{(X)},R^{(Y)}$ and stay below exchange limits $T_{X\to Y},T_{Y\to X}$ with $99\%$ probability. Expressed in standard form
$\min_{\bm{x},\bm{y}}$ $\bm{c^{T}}(\bm{x}+\bm{y})$ (20)
s.t. $0\leq\bm{x}+\bm{y}\leq 1$
[…]
with the first constraint ensuring each bid is procured at most once. To solve this problem, the chance-constraints have to be reformulated. For this, note that arbitrary weighted sums of gPC expansions result in a new gPC expansion
$\sum_{j=1}^{Z}z_{j}E_{j}(\bm{\xi})$ $=\sum_{j=1}^{Z}z_{j}\sum_{l=0}^{L}a_{l}^{j}\psi_{l}(\bm{\xi})$ (21)
$=\sum_{l=0}^{L}\bar{a}_{l}\psi_{l}(\bm{\xi})$
Figure 1: Illustration of the use case with four bids per zone. Two bids each are distributed across four locations with each two local random variables describing the temperature and irradiation.
where $\bar{a}_{l}=\sum_{j=1}^{Z}z_{j}a_{l}^{j}$. Using the properties in (4), the mean and variance of such a weighted sum can be expressed as functions of the weights
$\mu$ $=\sum_{j=1}^{Z}z_{j}a_{0}^{j}$ (22)
$\sigma$ $=\left\|\begin{bmatrix}a_{1}^{1}&\ldots&a_{1}^{Z}\\ \vdots&\ddots&\vdots\\ a_{L}^{1}&\ldots&a_{L}^{Z}\end{bmatrix}\begin{bmatrix}z_{1}\\ \vdots\\ z_{Z}\end{bmatrix}\right\|_{2}=\left\|(\bm{Az})_{Z}^{1}\right\|_{2}$
where $(\cdot)^{1}_{Z}$ indicates the index range of the matrix columns and vector. This allows a convex reformulation of the chance-constraints using quantile approximation, for instance
[…]
$\sum_{i=1}^{N_{X}+N_{Y}}x_{i}a_{0}^{i}+\lambda_{1}\left\|(\bm{Ax})^{1}_{N_{X}+N_{Y}}\right\|_{2}\geq R^{(X)}$
where $\lambda_{1}$ is the quantile factor for which $\mu+\lambda_{1}*\sigma$ corresponds to the $1^{st}$ percentile of the distribution. For normal distributions $\lambda_{1},\lambda_{99}=\mp 2.326$. Applying this reformulation to all chance-constraints in problem (20) results in a convex optimization problem.
$\min_{\bm{x},\bm{y}}$ $\bm{c^{T}}(\bm{x}+\bm{y})$ (24)
s.t. $0\leq\bm{x}+\bm{y}\leq 1$
$\sum_{i=1}^{N_{X}+N_{Y}}x_{i}a_{0}^{i}+\lambda_{1}\left\|(\bm{Ax})_{N_{X}+N_{Y}}^{1}\right\|_{2}\geq R^{(X)}$
$\sum_{i=1}^{N_{X}+N_{Y}}y_{i}a_{0}^{i}+\lambda_{1}\left\|(\bm{Ay})_{N_{X}+N_{Y}}^{1}\right\|_{2}\geq R^{(Y)}$
$\;\;\;\sum_{i=1}^{N_{X}}\;\;\;y_{i}a_{0}^{i}+\lambda_{99}\left\|(\bm{Ay})_{N_{X}}^{1}\right\|_{2}\leq T_{X\to Y}$
$\sum_{i=N_{X}+1}^{N_{X}+N_{Y}}x_{i}a_{0}^{i}+\lambda_{99}\left\|(\bm{Ax})_{N_{X}+N_{Y}}^{N_{X}+1}\right\|_{2}\leq T_{Y\to X}$
We solved a concrete example with $N_{X},N_{Y}=4$ bids per zone and $d=8$ random variables. In two situations. The bid functions are assumed to represent PV and HP capabilities. The random variables stand for the temperature and irradiation at four different locations with each two bids. Fig. 1 illustrates the setup.
III-1 Normal marginals
As $\lambda$ values are known exactly for normal distributions, the full process is first validated with arbitrary normally distributed marginals and linear bid functions. The correlation matrix $\bm{\Sigma}$ is fit from synthetic weather data of four different locations retrieved from renewables.ninja [9], however as there is no validation with real data, an arbitrary $\bm{\Sigma}$ could be used.
The polynomial basis is constructed from 9 monomials (all monomials of degree $\leq 1$). The quality of the gPC expansion is checked by drawing random samples from the Copula and evaluating both the bid functions and the expansion. The resulting distributions are displayed in Fig. 2. As constructed, the bid power distributions are indeed normally distributed. Moreover, the gPC expansion accurately approximates the bid functions. The maximum numerical error of the expansion was found to be on the order of $10^{-7}\%$
Figure 2: Sampled bid function distributions computed both with the bid functions and gPC expansions for the eight bids in the case of normally distributed marginals and linear bid functions.
Problem (24) is then solved with parameters $R^{(X)}$, $R^{(Y)}$, $T_{X\to Y}$, $T_{Y\to X}=100$ and $c_{1-8}=1$ to obtain optimal $\bm{x},\bm{y}$. The chance-constraints in (20) are validated by sampling both from the Copula and from the marginals in an uncorrelated. The resulting distributions are shown in Fig. 3. As expected, the optimization problem correctly satisfies the chance-constraints only for the correlated case. This results suggests, that the presented formalism works as intended.
Figure 3: Sampled procured power distributions computed with the gPC expansion in the case of normally distributed marginals. For comparison the PDFs both considering and not considering correlation were sampled. The empirical $1^{st}$ and $99^{th}$ percentiles are indicated for verification of the chance-constraints.
III-2 Arbitrary marginals
Obviously, bid functions are in general not linear in the random variables and the marginals not normally distributed. For example, solar irradiance uncertainty is frequently modelled using Beta distributions due to its bounded support. For validation in a more realistic setting, the bid functions are replaced by simple polynomial models of PV and HP capability and the marginals by representative Beta distributions.
The polynomial basis was constructed from 21 physically relevant monomials. The computation of the basis and all coefficients was achieved in roughly 2 minutes total with a non-parallelised implementation and $k=15$ quadrature points per dimension. Validating the gPC expansions with sampling resulted in the bid power distributions shown in Fig. 4. The maximum expansion error was found to be on the order of $10^{-6}\%$
Problem (24) was solved with parameters $R^{(X)}$, $R^{(Y)}=1000$, $T_{X\to Y}$, $T_{Y\to X}=500$ and $c_{1-8}=1$. The constraints validation by sampling resulted in the distributions shown in Fig. 5. Due to the non-linearity of the bid functions and the non-normal, non-symmetric marginals, the constraints with $\lambda$ values for normal distributions are no longer met perfectly.
In principle, for each chance-constraint a local search for optimal $\lambda$ values is possible. Another option is the use of a distributionally robust $\lambda$, which however leads to conservative estimates as $\lambda_{1}^{rob},\lambda_{99}^{rob}=\mp\sqrt{\frac{1-0.01}{0.01}}=\mp 9.950$
Figure 4: Sampled bid function distributions computed both with the bid functions and gPC expansions for the eight bids in the case of arbitrarily distributed marginals and polynomial bid functions.
Figure 5: Sampled procured power distributions computed with the gPC expansion in the case of arbitrarily distributed marginals. For comparison the PDFs both considering and not considering correlation were sampled. The empirical $1^{st}$ and $99^{th}$ percentiles are indicated for verification of the chance-constraints.
IV Conclusion
In this paper, we developed a tractable framework for modeling correlated stochastic energy resources using Polynomial Chaos Expansion combined with Gaussian Copulas. The approach enables to reformulate chance constraints on weighted sums of random functions as convex constraints. Several ideas were introduced to improve computational efficiency, including transformations for quadrature integration of Gaussian Copulas as well as precomputing expectation values for parallelization and effective dimensionality reduction.
The framework was validated on a numerical joint procurement problem with eight random variables and bids. First, using normally distributed marginals and linear bid functions, we showed that the reformulation correctly meets constraints for a correlated joint PDF. Then, for a more representative example with polynomial bid functions and Beta distributed marginals, a 21-dimensional orthonormal polynomial basis and corresponding PCE coefficients could be computed within two minutes. The optimization closely satisfies the chance constraints when using Gaussian quantile factors.
Key limitations include the reliance of dimensionality reduction techniques on the dimensionality of the random functions as well as the need for quantile factors $\lambda$ in the exact reformulation of chance constraints. While $\lambda$ is known for normal distributions and conservative robust bounds are available, its exact value remains generally unknown for arbitrary distributions.
Future work will focus on extending the framework to time-coupled problems, accounting for correlations across time steps. In addition, the formulation of joint chance constraints will be investigated.